OpenStudy (anonymous):

Julio purchased some power tools totaling $2,174 using a six-month deferred payment plan with an interest rate of 28.19%. He did not make any payments during the deferment period. What will the total cost of the power tools be if he must pay off the power tools within two years after the deferment period? $2,499.00 $3,297.84 $2,174.00 $4,122.30

OpenStudy (anonymous):

Start by using the compound interest formula.

OpenStudy (anonymous):

What would n be?

OpenStudy (anonymous):

I believe you have to do a cross product formula. 2174/x = 28.19/100

OpenStudy (anonymous):

I don't think so. If I have to use the compound interest formula it would be A=2174(1+0.2219/n)^n18 I just don't know what "n" would be.